OpenStudy (anonymous):

PLEASE HELP Find the inverse and determine whether or not its a function: f(x)=(8-2x)^2

OpenStudy (accessdenied):

We want to solve for x to find the inverse. f(x)= (8 - 2x)^2 sqrt( f(x) ) = |8 - 2x| +- sqrt( f(x) ) = 8 - 2x 8 +- sqrt( f(x) ) = -2x - (8 +- sqrt( f(x) ))/2 = x x = -(8 + sqrt( f(x) ))/2, and x = -(8 - sqrt( f(x) ))/2 f(x) = -(8 + sqrt(x)) / 2, and f(x) = -(8 - sqrt(x)) / 2 this inverse would not be a function (we could tell this from the start by horizontal line test on the original f(x), but we were asked to find the inverse anyways )

OpenStudy (anonymous):

So do i have to graph it in order to do the horizontal line test? I just won't have my graphing calculator with me for the test on this?

OpenStudy (accessdenied):

you would probably be asked to show the graph to use the horizontal line test, so yeah you'd just have to do it algebraically (like how I did it) to do it without graphing.

OpenStudy (anonymous):

well how do i know if its a function without using the test

OpenStudy (accessdenied):

basically, if you get more than one equation for x (with +-'s), any x-value will correspond to two different y values, which means that the inverse cannot be a function

OpenStudy (accessdenied):

Ex) y = x^2 +- sqrt(y) = x x = sqrt(y) and x = -sqrt(y) y = sqrt(x) and y = -sqrt(x) for any x, there's two y values.

OpenStudy (anonymous):

thank you makes sense!